OpenStudy (anonymous):

Find the x and y intercepts and graph the equation plotting the intercepts -4x+5y=-20 . please show work.

OpenStudy (annon):

for y intercept: 5y=4x-20 divid by 5: y=(4/5)x - 4 y intercept = -4

OpenStudy (annon):

for x intercept: -4x=-5y-20 multiply -1 to both sides: u get: 4x=5y+20 divid each side by 4: x=(5/4) y + 5 x intercept = 5

OpenStudy (annon):

get it?

OpenStudy (anonymous):

i guess just trying to graph